# TV-TREES: Multimodal Entailment Trees for Neuro-Symbolic Video Reasoning

### 01 Mar 2024 v2

@@ Abstract
It is challenging to perform question-answering over complex, multimodal content such as television clips. This is in part because current video-language models rely on single-modality reasoning, have lowered performance on long inputs, and lack interpretability. We propose TV-TREES, the first multimodal entailment tree generator. TV-TREES serves as an approach to video understanding that promotes interpretable joint-modality reasoning by producing trees of entailment relationships between simple premises directly entailed by the videos and higher-level conclusions. We then introduce the task of multimodal entailment tree generation to evaluate the reasoning quality of such methods. Our method’s experimental results on the challenging TVQA dataset demonstrate interpretable, state-of-the-art zero-shot performance on full video clips, illustrating a best of both worlds contrast to black-box methods.
@@ 1 Introduction
Figure 1: A (a) QA pair and (b) corresponding video clip and dialogue from the TVQA dataset Lei et al. (2018) and (c) a multimodal entailment tree, recursively produced by our approach (top-down). Trees are created through recursively retrieving atomic evidence from the transcript and video frames and decomposing the QA pair into compositionally equivalent hypotheses until each can be directly entailed by the retrieved evidence.
@@ 2.1 VideoQA
QA over images makes up a large portion of multimodal question-answering work Zou and Xie (2020). VideoQA benchmarks constitute a smaller portion of this area Zhong et al. (2022) and often focus on simple content and questions [CITE], but some recent videoQA datasets have targeted models’ commonsense knowledge and inference ability Lei et al. (2018); Zadeh et al. (2019). Recently, vision-and-language transformers have substantially improved performance on these videoQA tasks [CITE], and can often reason over complex content without an external knowledge base Kim et al. (2021); Wang et al. (2021b); Salin et al. (2022).
In contrast to these video-language models, Khurana and Deshpande (2021) highlight altenrative deep learning strategies for video QA such as attention-free methods, attention-based methods, memory network methods, and hierarchical reinforced methods. Notably, Zhao et al. (2018, 2020) propose a hierarchical encoder-decoder model that uses adaptive video segmentation based on the question contents. Related works consider graph networks for video understanding Wang et al. (2021a); Gu et al. (2021); Liu et al. (2022). While these models scale to longer videos more successfully, their performance suffers compared to transformer-based approaches.
@@ 3.2 Evaluation
To serve as a secondary and distinct objective from raw VideoQA performance, we propose an evaluation method for assessing the reasoning quality of multimodal entailment trees inspired by Weir et al.’s work on scoring compositional entailments Weir et al. (2024). Informal logic theory posits that natural language arguments may be evaluated in terms of their acceptability, relevance, and sufficiency Johnson and Blair (1977). We consider each node within an entailment tree as an “argument" and consider these qualia as guidelines for comprehensive entailment tree evaluation. Below, we formulate these three qualia through an information theoretic lens to establish a set of evaluation metrics. We use the Shannon definition of information gain,
@@ 4.1 System overview
The interaction of these three parts is illustrated in Algorithm 1. Given a hypothesis $h$, transcript sample $D^{\prime}\subseteq D$ and video sample $V^{\prime}\subseteq V$, the system first returns evidence from the transcript relevant to $h$ (line 1) and identifies whether any of it entails $h$ (2). If such evidence was retrieved, $e$ is set to the best sample (4) and the leaf node is returned (17). Otherwise, $h$ is decomposed into sub-hypotheses $h_{0}$ and $h_{1}$ (8) and the algorithm is recursively called on these newly constructed sub-problems (9-10), treating the generated sub-proofs as explanation $e$ (11). If textual evidence cannot be found for the current node nor any of the downstream nodes (14), then the visual evidence in sample $V^{\prime}$ is sampled, filtered, (13) and assigned to $e$ where applicable (15) in the same manner as the text content.
@@ 4.2 Preprocessing
Evidence Localization    Given the hypothesis, TV-TREES attempts to identify a temporal window to sample evidence from based on the dialogue. We use a cross-encoder model trained on the MS MARCO passage ranking task Bajaj et al. (2016) to rank six-line transcript passages on their computed similarity with the generated hypothesis. We use a sliding window to calculate scores for every potential sample and return the highest scoring excerpt. If a sufficient window is identified, the vision pipeline inherits this window. If no sufficient dialogue sample is found, the system uses all video frames as the evidence bank, omitting text entirely.
@@ 4.3 Evidence Retrieval
Existing natural language inference (NLI) models are not well-suited for classifying entailments within highly contextual and social dialogue, which often insinuate meaning not directly stated within the text. Instead of producing an entirely new dataset for the domain of dialogue NLI, we use GPT-3.5 to generate a set of natural language inferences about the dialogue sample written in the style as data points in a dataset akin to SNLI Bowman et al. (2015b), conditioned on a question form of the hypothesis, $q$. Presenting the question under discussion in the interrogative form significantly reduces the hallucination rate compared to passing in the original hypothesis. $q$ is also generated via GPT-3.5 taking the hypothesis $h$ as input.
@@ 4.4 Evidence Filtering
We use a cross-encoder trained on SNLI and MultiNLI to determine whether any of the retrieved evidence sufficiently entails the hypothesis. We accept any sample that achieves a logits score above a certain threshold for the "entailment" label.
@@ 4.6 Visual Reasoning
We pass in the questions generated in Section 5.3 alongside video frames from the localized evidence window (if applicable) sampled at 2 FPS into a vision-language model. In our experiments, we use LLaVA-7B Liu et al. (2023). To encourage conservative classifications, in addition to asking for “yes" and “no" answers we encourage the model to respond with “not enough information" if it is unsure or the image does not provide sufficient evidence. If more than 10% of the frames in the window result in an affirmative answer from the VLM model, the visual content is considered to contain sufficient entailing evidence and the frame with the highest logits score is returned. If no frames result in an affirmative answer, no appropriate visual evidence entails the hypothesis. The LLaVA-7B prompt is included alongside the GPT prompts in Appendix A.
We also use GPT-3.5 to anonymize the question generated in section 4.1, replacing character names with common nouns such as "person". We query LLaVA-7B on each frame individually, using the anonymized question as textual input. We compare the performance of this approach to providing the original question in Appendix B, but find that the modification makes marginal difference (approximately one-point lower performance on average).
@@ 5.1 Human Evaluations
Considering the three evaluation metrics described in Section 3.2 (acceptability, relevance, and sufficiency), we evaluate trees along these qualia through three annotation tasks. The first task provides annotators with the visual or text evidence assigned to the leaf nodes by the algorithm and ask them to assess the correctness of the leaf node hypotheses on a scale of 1-5 (acceptability) based on that evidence. The second task provides annotators with $(h_{0},h^{\prime})$ pairs from branching nodes and asks if the child hypothesis $h^{\prime}$ is relevant to the parent $h_{0}$ (relevance). The third task provides annotators with a full hypothesis triplet $(h_{0},h_{1},h_{2})$ from a branching node with parent $h_{0}$ and child premises $h_{1}$ and $h_{2}$ and asks (1) whether $h_{1}$ and $h_{2}$ each introduce distinct information (the other facet of relevance, we also call this distinctness for disambiguation purposes), and (2) if $h_{0}$ introduces information not provided by $h_{1}$ and $h_{2}$ together, to check for entailment (sufficiency). Through these tasks, annotators are also asked to indicate if any of the hypotheses or premises are malformed or otherwise uninterpretable (the other facet of acceptability).
@@ 5.2 GPT Evaluations
We take the qualia outlined in Section 3.2 and write three GPT-4 prompts for (1) correct leaves in the text domain, (2) correct leaves in the vision domain, and (3) the remaining three checklist items. Correctness check prompts are modality dependent as we use GPT-4V for vision evaluations, and separate both from the remaining checklist items as only the leaves must be evaluated for evidence-centric correctness. We use the same scoring values as in the human evaluations, and pass in twelve decompositions per prompt for the text prompts. These prompts are included in full in Appendix E.
@@ 6 Experiments
We evaluate TV-TREES on the TVQA dataset, comparing its performance against a text-only version of the architecture and competing zero-shot VideoQA approaches. We compare all approaches in terms of QA accuracy, and compare the entailment tree generation methods in terms of tree quality as described in Section 5.
@@ 6.1 Setup
We instantiate TV-TREES as it is described in Section 4, setting the maximum recursion depth to $k=2$, or allowing trees with up to 3 levels. Our experiments focus on the multiple choice VideoQA domain, and so we consider a question’s correct answer to be the answer that results in a complete tree. In the case that the system does not successfully complete any tree for the five answer candidates, we consider the answer candidate with the "most complete" tree to be the correct answer, breaking ties by the average entailment score at each node. When complete trees are generated for multiple answers, we break ties in the same way.
@@ 6.2 Evaluation on TVQA
Results   We report overall accuracy alongside qualitative comparisons between the approaches in Table 4. As shown in the table, TV-TREES outperforms existing zero-shot methods when using full clips, but still shows significant room for future improvements. Notably, the text-only model outperforms joint-modality methods, and the joint modality model only improves performance modestly, suggesting that the language modules of TV-TREES are more robust and performance could be further increased through improvements to the vision pipeline. This is further shown in the ablation experiment results in Table 3, which suggests that vision evidence alone allows TV-TREES to complete trees for only 19.7% of the questions compared to 51.5% and 69.5% for text-only and joint-modality models, respectively.
@@ 6.3 Proof Scoring
We evaluate these sampled trees using the automatic GPT4 approach as described in Section 5.2. We then sample 200 proofs from this set (evenly distributed across modalities and complexity) and we annotate this set with human annotators from Amazon Mechanical Turk as described in Section 5.1. For human annotations, we identify careful annotators through a preliminary pilot task where each annotator’s work is scored by hand, and only high-scoring annotators are invited to annotate the full proofs. More information regarding these crowdsourced annotations are included in Appendix C.
For scoring acceptability, we provide the scorer with the localized dialogue retrieved by the cross encoder model described in Section 4.2, and the video frames that achieved the highest logits scores during VQA inference, depending on the modality. We report results in Table 2.
@@
We find that, unsurprisingly, the majority of errors in the produced trees stem from acceptability issues. According to human evaluations, the visual module produces lower quality inferences than the textual modules do. This is not surprising, as we are able to include additional entailment filters for the textual reasoning steps to remove lower quality predictions before constructing the final entailment trees, whereas we do not have similar methods in place for visual inference. Based on these results, introducing stronger entailment classifiers for both domains may significantly improve performance on tree evaluation as well as on general VideoQA.
@@ 7 Conclusion
We introduce the first neuro-symbolic entailment tree generator for multimodal content to improve robustness, reliability, interpretability, and scalability of video-language understanding systems. We focus on the application of narrative-driven VideoQA, and show that our approach achieves state-of-the-art results on the zero-shot TVQA benchmark with full video clips. We also propose the task of multimodal entailment tree generation for the assessment of generated tree reasoning quality, establishing an information-theoretic evaluation method grounded in informal logic theory. Experimental results suggest that such interpretable, neuro-symbolic approaches to video understanding are a strong alternative to existing methods and present exciting directions for future research.

### 11 Mar 2024 v3

@@ Abstract
It is challenging to perform question-answering over complex, multimodal content such as television clips. This is in part because current video-language models rely on single-modality reasoning, have lowered performance on long inputs, and lack interpretability. We propose TV-TREES, the first multimodal entailment tree generator. TV-TREES serves as an approach to video understanding that promotes interpretable joint-modality reasoning by producing trees of entailment relationships between simple premises directly entailed by the videos and higher-level conclusions. We then introduce the task of multimodal entailment tree generation to evaluate the reasoning quality of such methods. Our method’s experimental results on the challenging TVQA dataset demonstrate interpretable, state-of-the-art zero-shot performance on full video clips, illustrating a best-of-both-worlds contrast to black-box methods.
@@ 1 Introduction
Figure 1: A (a) QA pair and (b) corresponding video clip and dialogue from the TVQA dataset Lei et al. (2018) and (c) a multimodal entailment tree, recursively produced by our approach (top-down). Trees are created by recursively retrieving atomic evidence from the transcript and video frames and decomposing the QA pair into compositionally equivalent hypotheses until each can be directly entailed by the retrieved evidence.
@@ 2.1 VideoQA
QA over images makes up a large portion of multimodal question-answering work Zou and Xie (2020). VideoQA benchmarks constitute a smaller portion of this area Zhong et al. (2022) and often focus on simple content and questions Jang et al. (2017), but some recent VideoQA datasets have targeted models’ commonsense knowledge and inference ability Lei et al. (2018); Zadeh et al. (2019). Recently, vision-and-language transformers have substantially improved performance on these VideoQA tasks Zhou et al. (2020), and can often reason over complex content without an external knowledge base Kim et al. (2021); Wang et al. (2021b); Salin et al. (2022).
In contrast to these video-language models, Khurana and Deshpande (2021) highlight alternative deep learning strategies for VideoQA such as attention-free methods, attention-based methods, memory network methods, and hierarchical reinforced methods. Notably, Zhao et al. (2018, 2020) propose a hierarchical encoder-decoder model that uses adaptive video segmentation based on the question contents. Related works consider graph networks for video understanding Wang et al. (2021a); Gu et al. (2021); Liu et al. (2022). While these models scale to longer videos more successfully, their performance suffers compared to transformer-based approaches.
@@ 3.2 Evaluation
To serve as a secondary and distinct objective from raw VideoQA performance, we propose an evaluation method for assessing the reasoning quality of multimodal entailment trees inspired by Weir et al.’s work on scoring compositional entailments Weir et al. (2024). Informal logic theory posits that natural language arguments may be evaluated in terms of their acceptability, relevance, and sufficiency Johnson and Blair (1977). We consider each node within an entailment tree as an “argument" and consider these qualia as guidelines for comprehensive entailment tree evaluation. Below, we formulate these three qualia through an information-theoretic lens to establish a set of evaluation metrics. We use the Shannon definition of information gain,
@@ 4.1 System overview
The interaction of these three parts is illustrated in Algorithm 1. Given a hypothesis $h$, transcript sample $D^{\prime}\subseteq D$, and video sample $V^{\prime}\subseteq V$, the system first returns evidence from the transcript relevant to $h$ (line 1) and identifies whether any of it entails $h$ (2). If such evidence was retrieved, $e$ is set to the best sample (4), and the leaf node is returned (17). Otherwise, $h$ is decomposed into sub-hypotheses $h_{0}$ and $h_{1}$ (8) and the algorithm is recursively called on these newly constructed sub-problems (9-10), treating the generated sub-proofs as explanation $e$ (11). If textual evidence cannot be found for the current node nor any of the downstream nodes (14), then the visual evidence in sample $V^{\prime}$ is sampled, filtered, (13) and assigned to $e$ where applicable (15) in the same manner as the text content.
@@ 4.2 Preprocessing
Evidence Localization    Given the hypothesis, TV-TREES attempts to identify a temporal window to sample evidence based on the dialogue. We use a cross-encoder model trained on the MS MARCO passage ranking task Bajaj et al. (2016) to rank six-line transcript passages on their computed similarity with the generated hypothesis. We use a sliding window to calculate scores for every potential sample and return the highest-scoring excerpt. If a sufficient window is identified, the vision pipeline inherits this window. If no sufficient dialogue sample is found, the system uses all video frames as the evidence bank, omitting text entirely.
@@ 4.3 Evidence Retrieval
Existing natural language inference (NLI) models are not well-suited for classifying entailments within highly contextual and social dialogue, which often insinuates meaning not directly stated within the text. Instead of producing an entirely new dataset for the domain of dialogue NLI, we use GPT-3.5 to generate a set of natural language inferences about the dialogue sample written in the style as data points in a dataset akin to SNLI Bowman et al. (2015b), conditioned on a question form of the hypothesis, $q$. Presenting the question under discussion in the interrogative form significantly reduces the hallucination rate compared to passing in the original hypothesis. $q$ is also generated via GPT-3.5 taking the hypothesis $h$ as input.
@@ 4.4 Evidence Filtering
We use a cross-encoder trained on SNLI and MultiNLI to determine whether any of the retrieved evidence sufficiently entails the hypothesis. We accept any sample that achieves a logit score above a certain threshold for the "entailment" label.
@@ 4.6 Visual Reasoning
We pass in the questions generated in Section 5.3 alongside video frames from the localized evidence window (if applicable) sampled at 2 FPS into a vision-language model. In our experiments, we use LLaVA-7B Liu et al. (2023). To encourage conservative classifications, in addition to asking for “yes" and “no" answers we encourage the model to respond with “not enough information" if it is unsure or the image does not provide sufficient evidence. If more than 10% of the frames in the window result in an affirmative answer from the VLM model, the visual content is considered to contain sufficient entailing evidence, and the frame with the highest logits score is returned. If no frames result in an affirmative answer, no appropriate visual evidence entails the hypothesis. The LLaVA-7B prompt is included alongside the GPT prompts in Appendix A.
We also use GPT-3.5 to anonymize the question generated in section 4.1, replacing character names with common nouns such as "person". We query LLaVA-7B on each frame individually, using the anonymized question as textual input. We compare the performance of this approach to providing the original question in Appendix B, but find that the modification makes a marginal difference (approximately one point lower performance on average).
@@ 5.1 Human Evaluations
Considering the three evaluation metrics described in Section 3.2 (acceptability, relevance, and sufficiency), we evaluate trees along these qualia through three annotation tasks. The first task provides annotators with the visual or text evidence assigned to the leaf nodes by the algorithm and asks them to assess the correctness of the leaf node hypotheses on a scale of 1-5 (acceptability) based on that evidence. The second task provides annotators with $(h_{0},h^{\prime})$ pairs from branching nodes and asks if the child hypothesis $h^{\prime}$ is relevant to the parent $h_{0}$ (relevance). The third task provides annotators with a full hypothesis triplet $(h_{0},h_{1},h_{2})$ from a branching node with parent $h_{0}$ and child premises $h_{1}$ and $h_{2}$ and asks (1) whether $h_{1}$ and $h_{2}$ each introduce distinct information (the other facet of relevance, we also call this distinctness for disambiguation purposes), and (2) if $h_{0}$ introduces information not provided by $h_{1}$ and $h_{2}$ together, to check for entailment (sufficiency). Through these tasks, annotators are also asked to indicate if any of the hypotheses or premises are malformed or otherwise uninterpretable (the other facet of acceptability).
@@ 5.2 GPT Evaluations
We take the qualia outlined in Section 3.2 and write three GPT-4 prompts for (1) correct leaves in the text domain, (2) correct leaves in the vision domain, and (3) the remaining three checklist items. Correctness check prompts are modality dependent as we use GPT-4V for vision evaluations, and separate both from the remaining checklist items as only the leaves must be evaluated for evidence-centric correctness. We use the same scoring values as in the human evaluations and pass in twelve decompositions per prompt for the text prompts. These prompts are included in full in Appendix E.
@@ 6 Experiments
We evaluate TV-TREES on the TVQA dataset, comparing its performance against a text-only version of the architecture and competing zero-shot VideoQA approaches. We compare all approaches in terms of QA accuracy and compare the entailment tree generation methods in terms of tree quality as described in Section 5.
@@ 6.1 Setup
We instantiate TV-TREES as it is described in Section 4, setting the maximum recursion depth to $k=2$, or allowing trees with up to 3 levels. Our experiments focus on the multiple-choice VideoQA domain, and so we consider a question’s correct answer to be the answer that results in a complete tree. In the case that the system does not successfully complete any tree for the five answer candidates, we consider the answer candidate with the "most complete" tree to be the correct answer, breaking ties by the average entailment score at each node. When complete trees are generated for multiple answers, we break ties in the same way.
@@ 6.2 Evaluation on TVQA
Results   We report overall accuracy alongside qualitative comparisons between the approaches in Table 4. As shown in the table, TV-TREES outperforms existing zero-shot methods when using full clips, but still shows significant room for future improvements. Notably, the text-only model outperforms joint-modality methods, and the joint-modality model only improves performance modestly, suggesting that the language modules of TV-TREES are more robust and performance could be further increased through improvements to the vision pipeline. This is further shown in the ablation experiment results in Table 3, which suggests that vision evidence alone allows TV-TREES to complete trees for only 19.7% of the questions compared to 51.5% and 69.5% for text-only and joint-modality models, respectively.
@@ 6.3 Proof Scoring
We evaluate these sampled trees using the automatic GPT4 approach as described in Section 5.2. We then sample 200 proofs from this set (evenly distributed across modalities and complexity) and we annotate this set with human annotators from Amazon Mechanical Turk as described in Section 5.1. For human annotations, we identify careful annotators through a preliminary pilot task where each annotator’s work is scored by hand, and only high-scoring annotators are invited to annotate the full proofs. More information regarding these crowdsourced annotations is included in Appendix C.
For scoring acceptability, we provide the scorer with the localized dialogue retrieved by the cross-encoder model described in Section 4.2, and the video frames that achieved the highest logits scores during VQA inference, depending on the modality. We report results in Table 2.
@@
We find that, unsurprisingly, the majority of errors in the produced trees stem from acceptability issues. According to human evaluations, the visual module produces lower-quality inferences than the textual modules do. This is not surprising, as we are able to include additional entailment filters for the textual reasoning steps to remove lower-quality predictions before constructing the final entailment trees, whereas we do not have similar methods in place for visual inference. Based on these results, introducing stronger entailment classifiers for both domains may significantly improve performance on tree evaluation as well as on general VideoQA.
@@ 7 Conclusion
We introduce the first neuro-symbolic entailment tree generator for multimodal content to improve the robustness, reliability, interpretability, and scalability of video-language understanding systems. We focus on the application of narrative-driven VideoQA and show that our approach achieves state-of-the-art results on the zero-shot TVQA benchmark with full video clips. We also propose the task of multimodal entailment tree generation for the assessment of generated tree reasoning quality, establishing an information-theoretic evaluation method grounded in informal logic theory. Experimental results suggest that such interpretable, neuro-symbolic approaches to video understanding are a strong alternative to existing methods and present exciting directions for future research.